# Training Language Models to Explain Their Own Computations

### v2

@@ 3 Privileged Access Improves Explanations
In this section, we (1) show that models can be fine-tuned to generate feature descriptions and generalize to new features and feature bases, and (2) provide evidence that this capability relies on privileged access. Our experiments focus on feature description (§2.2), where we train explainer models to verbalize descriptions of the target model’s features. We study additional tasks (patching, input ablation) in §5.
@@ 3.1 Models and Training Data
We study Llama-3.1-8B (grattafiori2024llama3herdmodels) as the target model, and test five different explainer models (Llama-3.1-8B, Llama-3-8B, Llama-3.1-8B-Instruct, Llama-3.1-70B, Qwen3-8B; yang2025qwen3technicalreport). [^5] Details can be found in Section A.1. We select the explainer models that differ from the target model in attributes such as architecture (Qwen3-8B), size (Llama-3.1-70B), and training regimen (Llama-3.1-8B-Instruct), to investigate the effect of each of these factors. We fine-tune each model to be an effective explainer as described in Equation˜3. For Llama-3.1-8B, we report results with both LoRA (hu2022lora) and full-parameter fine-tuning, while for Llama-3.1-70B we only report LoRA due to compute constraints, and for other models we only report full-parameter fine-tuning.
For models that require projection layers $\Pi_{\ell}$ due to hidden-size mismatch (Qwen3-8B, Llama-3.1-70B), we randomly initialize those layers and train them with the rest of the parameters. [^6] For LoRA training, these layers are included in the set of parameters trained via low-rank updates. We additionally train a version of Llama-3.1-70B where the projection layer is pre-trained to minimize $\|h^{\mathcal{E}}_{\ell,t}-\Pi_{\ell}\cdot h^{\mathcal{M}}_{\ell,t}\|_{2}$ on FineWeb samples (penedo2024the), or the distance between the the layer-$\ell$, token-$t$ explainer activation $h^{\mathcal{E}}_{\ell,t}$ , and the projected target activation $\Pi_{\ell}\cdot h^{\mathcal{M}}_{\ell,t}$ from the same layer, token location.
To generate training data, we use Llama-Scope (he2024llamascope) to obtain residual stream SAE features for Llama-3.1-8B target models, which we use for the direction $v$. We use the Llama-3.1-8B-LXR-32x features, which maps the residual stream after each layer to 131K features (32 times larger than the residual hidden dimension). [^7] We run additional experiments with Gemma 2 (gemmateam2024gemma2improvingopen), whose data generation details can be found in Section B.1. Ground-truth explanations $E$ for each feature were obtained from Neuronpedia (neuronpedia).
@@ 3.2 Evaluation Set and Metrics
We train on a subset of SAE features $S_{train}$, then evaluate generalization both in-distribution and out-of-distribution, on three types of features:
Held-out SAE features (SAE).
We hold out a subset of 1550 SAE features (50 per layer) for testing.
Full Activations (ACT).
We test out-of-distribution generalization to full residual stream activations $h_{(\ell,t)}(x)$ on inputs $x$ sampled from FineWeb (penedo2024the).
@@
Differences between counterfactual pairs of inputs allow us to isolate the effect of a single pointwise change. We create counterfactual pairs $(x,x^{\prime})$ from the Counterfact dataset (meng2022locating) using the procedure described in Section˜2.3 and extract activation differences $v=h_{\ell,t}(x)-h_{\ell,t}(x^{\prime})$.
On each of these evaluation sets, we consider two types of metrics:
LM judge.
For SAE features, gold descriptions are available, so we use an LM judge to assess the similarity of the predicted description to the gold description, on a scale of 0 (no similarity) to 1 (very similar), in increments of 0.25. We use GPT-4.1-mini as the judge; the exact prompt can be found in Section˜A.2. Human assessment of 200 samples of the LM judge results shows 81.25% human agreement. [^8] Details of this assessment can be found in Section A.2.
Simulator score.
We also use simulator score, described in Equation˜2, which measures the Pearson correlation between the model’s true activations and its predicted activations based on the explanation $\mathcal{E}(v)$ on each sample $x$.
@@ 3.3 Baselines
To evaluate whether explainer models exhibit nontrivial generalization to unseen features, we compare to a nearest-neighbor baseline that, given a new feature or representation, finds the single most similar feature from the training set, and outputs its associated description. Let $S_{\text{train}}$ denote the set of features used for training, and let $S_{\text{train},\ell}$ denote the subset of SAE features at layer $\ell$. We evaluate two nearest-neighbor baselines:
  1. 1.
    Layerwise nearest neighbor. We restrict retrieval to features from the same layer $\ell$, so that the retriever explanation is $\mathcal{E}_{\text{NN-layer}}(v,\ell)=D\!\left(\arg\max_{v_{i}\in S_{\text{train},\ell}}\langle v_{i},v\rangle\right)$, where $D$ is the function mapping training features to their explanations.
  2. 2.
    All-layer nearest neighbor. We retrieve from the training feature set across all layers: $\mathcal{E}_{\text{NN-all}}(v)=D\!\left(\arg\max_{v_{i}\in S_{\text{train}}}\langle v_{i},v\rangle\right)$.
SelfIE.
Following SelfIE (chen2024selfie), we directly patch $v$ into the embedding layer of a non-finetuned model, [^9] The original SelfIE paper patches representations into the third layer, but prior work has found that patching into the first few layers work about equivalently well (which we also find). Thus, we patch every activation into the embedding layer. and prompt the LM to define the feature. We use standard prompts found to work well for self-explanations (kharlapenko2024self). [^10] SelfIE prompts can be found in Section F.1. Prior works have found that SelfIE is sensitive to the scale of $v$ when it is inserted (kharlapenko2024self). We report results on the best explanation across 5 possible scales ($v$, $5v$, $10v$, $25v$, and $50v$) for each example, representing an upper bound on SelfIE performance.
We also procure gold SAE labels from Neuronpedia on the held-out SAE features. Note that the LM judge evaluates with respect to the same gold descriptions, meaning the gold SAE labels serve as a top-line number for the LM judge metric.
@@
NN-all 58.5±1.0 33.7±0.8 38.9±0.7 18.5±0.6
NN-layer 53.9±0.9 29.4±0.7 40.7±0.7 28.6±0.7
SelfIE Best of 5
    Llama-3.1-8B 40.1±1.0 36.4±0.8 43.3±0.6 21.0±0.6
    Llama-3.1-8B-Instr. 40.4±0.9 36.4±0.8 43.5±0.7 20.1±0.6
@@ 3.4 Results
We report results for our trained explainers, as well as each of the baselines, in Table˜1. Based on these results, we make several observations:
Trained explainers are effective.
For every feature type and metric, training Llama-3.1-8B to explain itself outperforms all of the baselines by a significant margin. It even outperforms the gold SAE labels when scored via the simulator, due to noise in the ground-truth feature labels themselves (see error analysis in Section˜B.5).
For explainers, matching the subject matters more than parameter count.
Among all trained models, the Llama-3.1-8B and Llama-3-8B models consistently performed the best across all four evaluations, while the Llama-3.1-8B-Instruct model was consistently a few (3-5) points behind. Meanwhile, the Qwen3-8B and the randomly-initialized Llama-3.1-70B model consistently performed the worst across all four evaluations, sometimes even underperforming baselines on the out-of-distribution evaluations. In this case, with Llama-3.1-8B as the target model, a Llama-3.1-8B explainer is significantly better than Llama-3.1-70B, showing that matching the target model matters more than a 9x increase in parameter count in this case.
Similar models are better explainers.
More generally, we observe that models closest in architecture and training regimen to Llama-3.1-8B perform the best, while models that deviate in one or both perform worse: Llama-3.1-8B and Llama-3-8B share architecture and deviate slightly in pre-training, while Llama-3.1-8B-Instruct has the same architecture but has an additional instruction-tuning step. Qwen3-8B and Llama-3.1-70B both deviate completely in either training regimen or architecture, with Qwen3-8B differing significantly in training (but having a similar architecture) and Llama-3.1-70B differing in architecture (but sharing similar pre-training regimen).
Alignment between activations improves verbalization capability.
To formalize the observation above, one way of quantifying the privileged access hypothesis is to posit that activation similarity between the explainer and target model predicts explainer performance. To investigate this, we compare the explanation capabilities of Llama-3.1-70b with both pre-trained projection and randomly initialized projection as a proxy for maximally and minimally aligned activation, respectively, given the model. As reported in Table˜1, we can recover a significant fraction of performance when we start training from the pre-trained projection, performing 15% better on SAE explanations, 2.7 times better on real activations, and 1.7 times better on activation differences.
Additionally, we quantify the activation similarity of (a subset of) models reported in Table˜1 with respect to the target model, and plot activation similarity vs. explainer performance. These plots can be found in Section˜B.2. Visually, they suggest that activation alignment is correlated with privileged access, in line with our findings above.
Taken together, these findings lead us to the following conclusion:
In feature description tasks, models’ privileged access to their internals improves explanations: they are better explained by fine-tuned versions of themselves than by other models.
Error Analysis.
Finally, we performed a qualitative error analysis of 100 samples where the LM judge scored the explainer’s prediction as 0 (details in Section˜B.5). We found that most errors stem from issues in the “gold” SAE labels from Neuronpedia (consistent with our finding that Llama-3.1-8B can outperform the gold labels), LM judge scoring inconsistencies, and inherent prediction difficulty from sparsely activating features. Only 27% of errors are attributed to genuine explainer failures.
@@ 4 Privileged Access Confers Data-Efficiency
To answer where and why privileged access may be useful, we investigate whether it confers data-efficiency advantages over alternatives. We train Llama-3.1-8B, Qwen3-8B, and the nearest neighbor baseline on subsets of the training data of various sizes and investigate scaling trends. The scaling curve for the held-out SAE features (LM judge evaluation) is reported in Figure˜5, while the scaling curves for the other metrics and evaluation settings can be found in Figure˜7.
Generally speaking, self-explanation is more data-efficient than either training another explainer model or nearest neighbors. At 0.8% of all training features (training on only 1024 samples per layer), Llama is already at 71% of its end-performance. Qwen is only at 35% of its end-performance and both forms of nearest neighbors (within- and across-layer) are at 55% of their end-performance. Note that the end-performance of Llama is already higher than both Qwen and nearest neighbors, meaning this gap is further amplified in the low data regime. Llama at 0.8% of training features achieves 92.9% of the end-performance of all-layer nearest neighbors and 77.4% of fully-trained Qwen. These results indicate that:
@@
This data-efficiency underscores the practical utility of self-explanation: annotating feature descriptions is expensive, and self-explanation training allows us to recover most of the performance with orders of magnitude less data. Figure˜5 shows that retrieving the top-1 SAE feature (i.e. nearest neighbors) from a feature dictionary of 32k samples per layer performs comparably to a self-explainer model trained on only 1k samples per layer—over a hundred times fewer features.
Finally, we speculate that the small amount of required data reflects what is minimally necessary to learn consistency between activation patterns and verbalizations: perhaps minimal training is necessary if models are simply reusing existing circuits that generate activation patterns for verbalization.
Figure 5: Training self-as-explainer is more data-efficient than alternatives. We plot scaling curves of each explainer: the number of training samples per layer is plotted against explanation quality, as judged by a LM judge on held-out SAE features. We find that matching the explainer to the target (Llama-3.1-8B) is generally the most data-efficient, while using another model as explainer (Qwen) or any form of nearest neighbor is much less effective in low-data regimes, taking over 1k-10k samples per layer to beat an untrained SelfIE baseline.
@@ 5 Generalization Across Tasks
To check whether self-explanation training and the privileged access hypothesis can be extended to other domains, we run preliminary experiments for two additional explanation types: predicting the outcomes of activation patching (§2.3) and input ablation (§2.4).
We examine models’ ability to explain results from activation patching experiments, which perturbs various parts of model components and examines which cause a change to the model output. This method is described in detail in Section˜2.3. We train two different explainer models—Llama-3.1-8B and Qwen3-8B—to explain each other as the target models, creating four different combinations of evaluation settings.
Training Data.
We use the CounterFact dataset (meng2022locating) as a source of counterfactual input pairs $x,x^{\prime}$, which consists of a corpus of factual subject–relation–object sentences like Paris (subject) is the capital of (relation) France (object). We prompt the model with the subject and relation and ask the model to predict the object . We then identify a counterfactual sentence $x^{\prime}$ with the same relation but differing subject and object (Rome is the capital of) and measure whether the most likely LM prediction changes (e.g. to Italy). We perform sampling to ensure an even split between changed and unchanged predictions, at each token position and layer chunk. See Section˜E.2 for details.
Baselines.
To measure the effect of fine-tuning, we prompt the pre-trained Qwen3-8B and Llama-3.1-8B models (without any explanation training) to generate explanations of the form described in Section˜2.3. The full prompts for zero-shot activation patching can be found in Section˜F.2. This is the analogous setting to the SelfIE baseline for feature description.
@@ 5.1 Activation Patching: Training LMs to Predict the Outcomes of Activation Interventions
Recall from Section˜2.3 that the explainer output typically includes two parts: whether the target output would changed under intervention (changes / remains unchanged), and the content of the change (to France). Based on this, we report three different types of metrics:
@@
    Has-Changed F1: Measures correctness of the first part of the explainer’s prediction, of whether the target’s output would change under intervention. We report the Macro F1 scores over the “changed” and “unchanged” classes.
@@
    Content Match: Measures correctness of the second part of the explainer’s prediction, of whether the explainer correctly predicts the content of target model’s prediction under intervention.
@@
Target Explainer Exact Match Has-Changed F1 Content Match
Qwen3-8B Qwen3-8B 64.0±0.4 80.2±0.3 71.0±0.4
– activation 59.9±0.4 76.5±0.4 68.6±0.4
– layer 65.0±0.4 80.0±0.3 72.7±0.4
– token 63.9±0.4 79.4±0.3 71.7±0.4
Llama-3.1-8B 54.1±0.4 73.9±0.4 65.4±0.4
Qwen-3-8B (Untrained) 5.02±0.2 24.4±0.8 18.78±0.3
Llama-3.1-8B Llama-3.1-8B 48.6±0.7 77.5±0.6 54.6±0.7
– activation 45.2±0.7 73.9±0.6 50.8±0.7
– layer 49.7±0.7 76.9±0.6 55.7±0.7
– token 47.3±0.7 76.1±0.6 53.4±0.7
Qwen3-8B 41.7±0.7 75.1±0.6 47.6±0.7
Llama-3.1-8B (Untrained) 3.2±0.1 31.4±0.4 7.0±0.2
Table 2: A target model’s activation patching outcomes are best predicted by itself. We train Qwen3-8B and Llama-3.1-8B to verbalize their own activation patching outcomes, and compare against training the other model to explain their outcomes, and also against untrained versions of themselves. Scores (mean $\pm$ standard error) are reported for three metrics—exact match, has-changed F1, and content prediction. For each target model and metric, bold indicates no significant difference from the best entry (paired $t$-test, $p\geq 0.05$). We also conduct experiments ablating the activation, layer, and token information from the explainer’s input. Ablating layer and token each have neglible effect, potentially because they can already be inferred from the activation.
@@ Results
for activation patching can be found in Table˜2. We see that the privileged access hypothesis is supported: Llama-3.1-8B is best at explaining Llama-3.1-8B, while Qwen3-8B is best at explaining Qwen3-8B. All trained explainers perform better than their untrained versions, which generally guess the has-changed value randomly.
Ablations.
In addition to training each explainer model on each target model, we perform a set of ablations to isolate which component(s) of the input the explainer learned to condition on. The explainer input consists of three components: the activation $v$ to insert (– activations), the layers $\ell_{1:i}$ (– layer), and the token $t$ of the target model to patch into (– token). We train the explainer model with each aspect of the input ablated (conditioning only on the other two components) to generate the ground-truth explanations. [^11] Ablated prompts can be found in Section F.3. Our ablation studies suggest found that performance is generally robust to ablating the layer and token annotations in the input prompt. We believe this is because layer and token information is generally recoverable from the activation and input context alone. As evidence for this, we train Llama-3.1-8B to decode the layer and token from which an activation comes from, and get 98.7% accuracy (see Section˜C.1).

### v3

@@ 3 Privileged Access Improves Explanations
In this section, we show that (1) models can be fine-tuned to generate feature descriptions and generalize to new features and feature bases, and (2) this capability relies on privileged access. We focus on feature description (§2.2), and study additional tasks (patching, input ablation) in §5.
@@ 3.1 Models and Training Data
Models.
We use Llama-3.1-8B (Grattafiori et al., 2024) as the target model, [^5] We run additional experiments with Gemma 2 (Gemma Team et al., 2024) as target model, see Section C.1. and train five different explainer models (Llama-3.1-8B, Llama-3-8B, Llama-3.1-8B-Instruct, Llama-3.1-70B, Qwen3-8B; Yang et al., 2025). [^6] Details can be found in Section B.1. For Llama-3.1-8B, we report results with both LoRA (Hu et al., 2022) and full-parameter fine-tuning. For Llama-3.1-70B we only report LoRA; for other models we only report full fine-tuning.
Initializing $\Pi_{\ell}$.
For models that require projection layers $\Pi_{\ell}$ due to hidden-size mismatch (Qwen3-8B, Llama-3.1-70B), we randomly initialize $\Pi_{\ell}$ and train them with the rest of the parameters. [^7] For LoRA training, these layers are included in the set of parameters trained via low-rank updates. For Llama-3.1-70B, we also include results where $\Pi_{\ell}$ was pre-trained on $\min_{\Pi_{\ell}}\|h^{\mathcal{E}}_{\ell,t}-\Pi_{\ell}\cdot h^{\mathcal{M}}_{\ell,t}\|_{2}$ (distance between explainer and projected target activations) on FineWeb samples (Penedo et al., 2024),
Data.
We train on residual stream SAE features of Llama-3.1-8B target models from Llama-Scope (He et al., 2024). We use the Llama-3.1-8B-LXR-32x features, which maps the residual stream after each layer to 131K features (32 times larger than the residual hidden dimension). Ground-truth explanations $E$ for each feature were obtained from Neuronpedia (Lin, 2023).
@@ 3.2 Evaluation Set and Metrics
We train on a subset of SAE features, then evaluate generalization in- and out-of-distribution on three types of features:
Held-out SAE features (SAE): We hold out a subset of 1550 SAE features (50 per layer) for testing.
Full Activations (ACT): We test out-of-distribution generalization to full residual stream activations $h_{(\ell,t)}(x)$ on inputs $x$ sampled from FineWeb (Penedo et al., 2024).
@@
To measure explainer performance on each type of feature, we use two metrics. (1) For SAE features, gold descriptions exist, so an LM judge can assess the similarity of predictions to the gold description, on a scale of 0 to 1 in increments of 0.25. The LM judge has 81.25% human agreement pe Appendix˜A. (2) We also use simulator score (Equation˜2), measuring the Pearson correlation between true and predicted activations based on $\mathcal{E}(v)$ for each sample $x$.
@@ 3.3 Baselines
To evaluate whether explainer models exhibit nontrivial generalization to unseen features, we compare to a nearest-neighbor baseline that, given a new feature or representation, finds the single most similar feature from the training set, and outputs its associated description:
$\mathcal{E}_{\text{NN}}(v)=D\!\left(\operatorname*{arg\,max}_{v_{i}\in S_{\text{train}}}\langle v_{i},v\rangle\right).$
where $S_{\text{train}}$ denotes the set of features used for training. This gives us an estimate of roughly how well the explainer can do if it simply learned to memorize and interpolate samples from the training data. [^8] We also perform layer-wise nearest neighbors in Section B.2.
SAE
@@
Nearest Neighbors 58.5±1.0 33.7±0.8 38.9±0.7 18.5±0.6
SelfIE Best of 5 40.1±1.0 36.4±0.8 43.3±0.6 21.0±0.6
@@
SelfIE.
Following SelfIE (Chen et al., 2024a), we directly patch $v$ into the embedding layer of a non-finetuned Llama-3.1-8B model, [^9] The original SelfIE paper patches representations into the third layer, but prior work has found that patching into the first few layers work about equivalently well (which we also find). Thus, we patch every activation into the embedding layer. and prompt it to define the feature. We use standard prompts found to work well for self-explanations (Kharlapenko et al., 2024). [^10] SelfIE prompts can be found in Section H.1. Because SelfIE is sensitive to the scale of $v$ when it is inserted (Kharlapenko et al., 2024), we report results on the best explanation across 5 possible scales ($v$, $5v$, $10v$, $25v$, and $50v$), representing an upper bound on SelfIE performance.
Gold SAE labels.
For held-out SAE features, we evaluate their gold Neuronpedia labels with the simulator.
@@ 3.4 Results
Results can be found in Table˜1. We observe:
Trained explainers are effective. For every feature type and metric, training Llama-3.1-8B to explain itself outperforms all baselines by a significant margin. This even includes gold SAE labels due to noise in the ground-truth feature labels themselves (see error analysis in Section˜A.3).
Alignment between activations improves verbalization capability. Among all trained models, the Llama-3.1-8B and Llama-3-8B models consistently outperform Llama-3.1-8B-Instruct by a few points and significantly outperform Qwen3-8B and Llama-3.1-70B, despite Llama-3.1-70B’s larger size. We posit that activation similarity between the explainer and target model predicts explainer performance.
We confirm this correlation visually by plotting explainer-target activation similarity vs. explainer performance in  Section˜C.2. the explanation capabilities of Llama-3.1-70b with the pre-trained projection and randomly initialized projection serve as a proxy for maximally and minimally aligned activation, respectively, given the model. As reported in Table˜1, we can recover a significant fraction of performance when we start training from the pre-trained projection, performing 14% better on SAE explanations, 2.7 times better on real activations, and 1.7 times better on activation differences. We conclude that:
For feature descriptions, privileged access improves model explanations: models are better explained by fine-tuned versions of themselves than by other models.
Figure 5: Training self-as-explainer is more data-efficient than alternatives. Scaling curves of training samples per layer vs. explanation quality (as measured either by LM judge or simulator) on three types of features (SAE, full activations, activation differences) show that matching explainer to target (Llama-3.1-8B) is more data-efficient than using a different model (Qwen) or nearest neighbors, especially in the low-data regime. The latter requires more samples per layer to outperform an untrained SelfIE baseline. Note that we aggregate 5 attempts of SelfIE with different strength, so the graph shows an upper bound of SelfIE performance.
@@ 4 Privileged Access Confers Data-Efficiency
To answer where and why privileged access may be useful, we investigate whether it confers data-efficiency advantages over alternatives. We train Llama-3.1-8B, Qwen3-8B, and the nearest neighbor baseline on subsets of the training data of various sizes and plot scaling curves against SelfIE baseline in Figure˜5.
[…]
@@
This data-efficiency underscores the practical utility of self-explanation: annotating feature descriptions is expensive, and self-explanation training allows us to recover most of the performance of other explanation techniques with orders of magnitude less data.
@@ 5 Generalization Across Tasks
We check if self-explanation and privileged access generalize to other domains by training models to explain outcomes of activation patching (§2.3) and input ablation (§2.4).
Models.
Across both tasks, we train two explainer models, Qwen3-8B and Llama-3.1-8B, to explain Qwen3-8B. We also report results with Llama-3.1-8B as the target model for activation patching and input ablations in Sections˜D.1 and E.1 respectively.
@@
Recall from Sections˜2.3 and 2.4 that the explainer output for both activation patching and input ablations typically includes two parts: whether the target output would changed under intervention (changes / remains unchanged), and the content of the change (to France). Based on this, we report three different types of metrics:
@@
    Has-Changed F1: Correctness when predicting whether the target’s output would change under intervention. We report the Macro F1 scores over the “changed” and “unchanged” classes.
@@
    Content Match: Correctness of predictions about the content of target model output under intervention.
@@
Baselines.
To measure the effect of fine-tuning, we prompt Qwen3-8B (without any explanation training) to generate explanations of its own activation patching and input ablation procedures. The full prompts for untrained activation patching are in Section˜H.2. and for untrained input ablations are in Section˜H.3. These baselines are analogous to the SelfIE baseline in Section˜3.3.
Activation Patching Input Ablation
Target Explainer Exact Match Has-Changed F1 Content Match Exact Match Has-Changed F1 Content Match
Qwen3-8B Qwen3-8B 64.0±0.4 80.2±0.3 71.0±0.4 83.4±1.0 87.0±1.1 90.6±0.8
Llama-3.1-8B 54.1±0.4 73.9±0.4 65.4±0.4 58.1±1.3 70.5±0.0 71.6±1.2
Qwen3-8B (Untrained) 5.02±0.2 24.4±0.8 18.8±0.3 8.9±1.5 44.4±1.6 35.3±2.5
Table 2: A target model’s intervention outcomes are best predicted by itself, shown through activation patching and input ablation. We train Qwen3-8B to verbalize its own intervention outcomes, and compare against training a different model (Llama-3.1-8B), and also against an untrained version of Qwen3-8B. Scores (mean $\pm$ standard error) are reported for three metrics—exact match, has-changed F1, and content prediction. For each task and metric, bold indicates no significant difference from the best entry (paired $t$-test, $p\geq 0.05$). Results with Llama-3.1-8B as the target model can be found in Tables˜5 and 6.
We train models to explain outcomes of activation patching, which perturbs a model component and examines whether it changes the model output, described in detail in Section˜2.3.
Training Data.
We use the CounterFact dataset (Meng et al., 2022) as a source of counterfactual input pairs $x,x^{\prime}$, which consists of a corpus of factual subject–relation–object sentences like Paris (subject) is the capital of (relation) France (object). We prompt the model with the subject and relation and ask the model to predict the object . We then identify a counterfactual sentence $x^{\prime}$ with the same relation but differing subject and object (Rome is the capital of) and measure whether the most likely LM prediction changes (e.g. to Italy). We perform sampling to ensure an even split between changed and unchanged predictions, at each token position and layer chunk. See Section˜G.2 for details.
@@ Results
for activation patching can be found in Table˜2. The privileged access hypothesis is supported: Qwen3-8B is best at explaining Qwen3-8B. All trained explainers perform better than their untrained versions, which generally guess the has-changed value randomly. We also include a set of ablation experiments in Section˜D.2 to isolate which input component(s) the explainer learned to condition on.